OpenStudy (anonymous):

collect the like terms 3/7r+3/7s-1/7r+16/7s

OpenStudy (anonymous):

rearranged that would be: =(3/7)r - (1/7)r + (3/7)s + (16/7)s =(2/7)r + (19/7)s